# Majorana edge and end states in planar Josephson junctions

### v2

@@ Abstract
We theoretically investigate the localization properties of Majorana states (MSs) in proximitized, planar Josephson Junctions (JJs) oriented along different crystallographic orientations and in the presence of an in-plane magnetic field and Rashba and Dresselhaus spin-orbit couplings. We show that two types of MSs may emerge when the junction transits into the topological superconducting state. In one case, referred to as end-like MSs, the Majorana quasiparticles are mainly localized inside the normal region at the opposite ends of the junction. In contrast, edge-like MSs extend along the opposite edges of the system, perpendicular to the junction channel. We show how the MSs can transit from end-like to edge-like and vice versa by tuning the magnetic field strength and/or the superconducting phase difference across the junction. In the case of phase-unbiased JJs the transition may occur as the ground state phase difference self-adjusts its value when the Zeeman field is varied.
@@ I Introduction
Majorana states (MSs) are zero-energy quasiparticle excitations predicted to appear localized at the boundaries of topological superconductors (TSCs) Kitaev (2001, 2003); Qi and Zhang (2011); Leijnse and Flensberg (2012); Beenakker (2013); Aguado (2017). The MSs obey a non-Abelian exchange statistic, which makes them promising candidates for realizing robust qubits with potential applications in fault-tolerant quantum computing Ivanov (2001); Nayak et al. (2008); Alicea et al. (2011); Aasen et al. (2016).
@@
Topological superconductivity (TS) can be engineered by using semiconductor nanowires with large spin-orbit coupling (SOC) and proximitized by s-wave superconductors Oreg et al. (2010); Sau et al. (2010a, b); Lutchyn et al. (2010); Rokhinson et al. (2012); Pientka et al. (2012); Mourik et al. (2012); Das et al. (2012); Deng et al. (2012, 2016); Manna et al. (2020), proximitized systems exposed to magnetic textures Klinovaja et al. (2012); Kjaergaard et al. (2012); Fatin et al. (2016); Matos-Abiague et al. (2017); Mohanta et al. (2019); Zhou et al. (2019); Marra and Cuoco (2017); Desjardins et al. (2019); Steffensen et al. (2021); Marra (2022); Güngördü and Kovalev (2022), and magnetic chains on s-wave superconductors Choy et al. (2011); Martin and Morpurgo (2012); Pientka et al. (2013); Nadj-Perge et al. (2013, 2014); Pawlak et al. (2016). Proximitized planar JJs have recently emerged as promising platforms for creating and manipulating MSs Kontos et al. (2002); Pientka et al. (2017); Pekker et al. (2013); Hell et al. (2017); Setiawan et al. (2019a, b); Yokoyama et al. (2014); Fornieri et al. (2019); Ren et al. (2019); Dartiailh et al. (2021); Hart et al. (2014); Laeven et al. (2020); Svetogorov et al. (2021); Lesser et al. (2021); Zhou et al. (2022); Stenger et al. (2019); Cayao et al. (2017); Scharf et al. (2019a); Zhou et al. (2020); Zhang et al. (2020); Woods and Stanescu (2020); Paudel et al. (2021); Salimian et al. (2021); Setiawan et al. (2017); Setiawan and Hofmann (2022); Banerjee et al. (2023); Pekerten et al. (2024a). In addition to the experimental advances in building such structures, proximitized planar JJs have also been shown to possess an enhanced parameter space supporting the topological superconducting state Pientka et al. (2017); Pekerten et al. (2024b).
@@
Magnetic and crystalline anisotropic effects have been predicted to appear in the Josephson junction with noncentrosymmetric materials Pakizer et al. (2021); Pekerten et al. (2022); Tanaka and Kashiwaya (1997); Tanaka et al. (2010). In particular, it has been shown that in the presence of SOC the Zeeman interaction yields a strong dependence of the system properties on the magnetic field direction. Furthermore, in systems with Rashba Bychkov and Rashba (1984) and Dresselhaus Dresselhaus (1955) SOCs the crystallographic orientation can affect the topological superconducting state, its robustness, and signatures Scharf et al. (2019b); Pakizer et al. (2021); Pekerten et al. (2022).
@@
Most previous investigations of TS in planar JJs have focused on end MSs, i.e., MSs that localize at the opposite ends of the junction with short localization lengths both along the junction ($\hat{y}$ direction, as shown in Fig. 1) and along the system edges perpendicular to the junction ($\hat{x}$ direction, as shown in Fig. 1) Pekerten et al. (2022); Zhou et al. (2020); Pakizer et al. (2021); Luethi et al. (2023); Zhou et al. (2022). However, theoretical evidence of the existence of edge MSs (i.e., MSs that are localized along the junction direction but spread along the entire system edges perpendicular to the junction) has been provided in previous works Sau et al. (2010c); Fornieri et al. (2019); Pakizer et al. (2021). The formation of MSs exhibiting anomalous multilocality in three-terminal Josephson junctions has also been proposed Nagae et al. (2024). Besides JJs, Majorana edge states can also emerge in hybrid superconductor/ferromagnet structures with helical magnetic textures Wu et al. (2017). Moreover, two-dimensional structures, typically associated with quantum anomalous Hall systems coupled to superconductors, have been shown to support chiral Majorana edge states Lian et al. (2018); Ménard et al. (2019); Beenakker and Oriekhov (2020) flowing around the sample edges in opposite directions and Majorana corner modes localized at vertices Zhu (2018); Liu et al. (2024); Winblad and Chen (2024).
@@ II Theoretical Model
Figure 1: (a) Schematic of a JJ consisting of a noncentrosymmetric semiconductor 2DEG (blue) in contact with two superconducting (S) leads (green). The $\hat{x}$ and $\hat{y}$ axes define the coordinate system in the junction’s reference frame. A top gate (not shown) over the normal (N) region can be used to modulate the Rashba SOC strength Dartiailh et al. (2021); Mayer et al. (2020). (b) Relevant angles in the junction coordinate system: $\varphi_{B}$ defines the direction of the in-plane magnetic field ($\mathbf{B}$) with respect to the $\hat{x}$ axis, while $\theta_{c}$ determines the orientation of the junction reference frame with respect to the semiconductor’s [100] crystallographic axis.
@@ II.1 Topological charge
where ${\rm Pf}\{...\}$ denotes the Pfaffian Schnyder et al. (2008); Ryu et al. (2010); Ghosh et al. (2010); Tewari and Sau (2012). The topological charge determines whether a system belonging to the D class is in a trivial ($Q=1$) or topological ($Q=-1$) phase Rieder et al. (2013); Adagideli et al. (2014); Pekerten et al. (2017); Stanescu et al. (2011); Stanescu and Tewari (2013); Pekerten et al. (2019).
It is worth noting that under some conditions determined by the SOC field, the magnetic field direction, and the junction crystallographic orientation, symmetric junctions may effectively belong to the BDI class Pakizer et al. (2021). In such cases, the topological phases are characterized by the $\mathbb{Z}$ topological invariant of the BDI class. Since the topological charge $Q$ is determined by the parity of the $\mathbb{Z}$ index, the topologically non-trivial regions of the BDI class consist of non-trivial D-class regions (composed of odd $\mathbb{Z}$ index subregions) enhanced by regions with even $\mathbb{Z}$ index Pientka et al. (2017); Setiawan et al. (2019a). We found that the topological gap in BDI class regions with multiple pairs of MSs is relatively small in the systems considered here. Therefore, our investigation focuses on regions that support only a single pair of MSs. The extent of these regions can be determined by examining how the topological charge $Q$ depends on the system parameters.
@@ II.2 Topological gap
Here $E_{0}$ and $E_{1}$ are the two lowest-energy states on the positive branch of the energy spectrum, respectively. Due to finite-size effects, the MSs localized at opposite ends (or edges) may overlap, so their energy (±$\pm E_{0}$) may slightly deviate from zero. Note that Eq. (8) can only be interpreted as the topological gap when the system is in the TS state.
@@
The magnitude of $\Delta_{top}$ has been shown to strongly depend on the junction’s crystallographic orientation $(\theta_{c})$, the spin-orbit angle $(\theta_{so})$ and the in-plane magnetic field orientation $(\varphi_{B})$, being optimal when the following relation is fulfilled Pakizer et al. (2021); Pekerten et al. (2022),
@@
As briefly discussed in the Introduction, end-like MSs are localized at the opposite ends of the junction channel, with short localization lengths both along the junction and along the system’s edges perpendicular to it. In contrast, edge-like MSs are localized solely along the junction direction but extend across the full length of the system’s edges perpendicular to the junction. We want to emphasize that the edge-like MSs examined here differ from both chiral Majorana states commonly associated with quantum anomalous Hall systems coupled to superconductors Lian et al. (2018); Ménard et al. (2019); Beenakker and Oriekhov (2020) and unidirectional Majorana edge states in noncentrosymmetric superconductors Daido and Yanase (2017).
To understand the origin of the extended nature of edge-like MSs, consider a simplified model of a junction sufficiently long such that the overlap between Majorana states at opposite edges is negligible. In this scenario, we can focus solely on one Majorana state, for instance, the one localized at the bottom edge ($y=0$). Assuming the S regions are infinitely wide ($W_{S}\to\infty$), the scattering states in the left S region can be expressed as Pakizer et al. (2021); Scharf et al. (2019b),
@@
where $\kappa$ describes the localization of the Majorana state along the junction $y$-direction, the subindexes $e$ and $h$ refer to electron-like and hole-like states, respectively, $s=\pm$ characterizes the spin, and $q_{e,s}$ ($q_{h,s}$) is the wave vector of the electron-like (hole-like) state with spinor $\chi_{e,s}$ ($\chi_{h,s}$) and spin $s$.
@@
The overall localization length of the Majorana state along the edge is given by the largest of the $l_{j,s}$. Hence, at $\phi=0$, end-like Majorana states appear when all wave vectors are complex. Conversely, if at least one wave vector is purely real, the corresponding scattering mode represents a propagating wave, leading to an edge-like MS.
@@
are satisfied, all coefficients of the quartic equation become real. This allows for purely real wave vectors, enabling the emergence of edge-like MSs if the system parameters are appropriately tuned. All the configurations considered in the numerical analysis presented in Sec. III are chosen in such a way that Eq. (14) and the condition for optimal topological gap [Eq. (9) are simultaneously fulfilled.
@@
$\tilde{\Delta}=\dfrac{(1-Q)\Delta_{top}}{2\Delta_{0}}\,{\rm sgn}(\xi)$ (15)
where $\xi$ s a parameter that characterizes the nature of the MSs, taking positive values for edge-like MSs and negative values for end-like MSs.
@@
$\tilde{\Delta}=\begin{cases}+\Delta_{top}/\Delta_{0}&\text{for D-class TS withedge-like MSs}\\ 0&\text{for trivial phase with no MSs}\\ 0&\text{for even BDI-class TS phase}\\ -\Delta_{top}/\Delta_{0}&\text{for D-Class with end-like MSs}\end{cases}.$ (16)
@@
For the junctions studied here, the probability density of edge-like MSs displays an oscillatory behavior along the edge, with multiple local maxima of comparable amplitudes. In contrast, the probability density oscillations of end-like MSs decay rapidly, featuring a large maximum at the middle of the junction edge, followed by a few significantly smaller maxima. Therefore, the localization nature of the MSs can be captured by the parametrization $\xi=n-n_{0}-1/2$, where $n$ is the number of probability density local maxima with comparable amplitudes. Since the typical number of local maxima for end-like MSs is smaller than 3, we set $n_{0}=3$ in the numerical simulations of Eq.(16) discussed below. For higher chemical potentials and/or larger $W_{N}$ and $W_{S}$ one may need to increase $n_{0}$ accordingly.
@@ III Results
We consider two types of junctions: (i) Al/HgTe Josephson junctions (JJs), where Rashba spin-orbit coupling (SOC) is the dominant effect, and (ii) Al/InSb JJs, where both Rashba and Dresselhaus SOC may be significant. The system parameters used in our calculations are provided in Appendix A. The numerical simulations were carried out by discretizing Eq. (1) on a mesh with a lattice constant of $a=10$ nm. Using the finite-difference approximation, we constructed the tight-binding (TB) form of the BdG Hamiltonian with the Kwant package Groth et al. (2014). The energy spectrum and wave functions were obtained by numerically diagonalizing the TB BdG Hamiltonian on a finite lattice. The energy spectrum was then used to compute the topological gap [Eq. (8)], the phase-dependent part of the free energy [Eq. (16)], and the ground-state phase [Eq. (16)]. The topological charge [see Eq. (7)] was computed by using the system TB BdG Hamiltonian with imposed translational invariance along the junction direction.

### v3

@@ Abstract
We theoretically investigate the localization properties of Majorana states (MSs) in proximitized, planar Josephson Junctions (JJs) oriented along different crystallographic orientations and in the presence of an in-plane magnetic field and Rashba and Dresselhaus spin-orbit couplings. We show that two types of MSs may emerge when the junction transits into the topological superconducting state. In one case, referred to as end-like MSs, the Majorana quasiparticles are mainly localized inside the normal region at the opposite ends of the junction. In contrast, edge-like MSs extend along the opposite edges of the system, perpendicular to the junction channel. We show how the MSs can transit from end-like to edge-like and vice versa by tuning the magnetic field strength and/or the superconducting phase difference across the junction. In the case of phase-unbiased JJs the transition may occur as the ground state phase difference self-adjusts its value when the Zeeman field is varied. We propose exploiting the extended nature of edge-like MSs as effective interconnects enabling the coupling between topological states in adjacent planar JJs. The impact of electrostatic disorder on the MSs is also analyzed.
@@ I Introduction
Majorana states (MSs) are zero-energy quasiparticle excitations predicted to appear localized at the boundaries of topological superconductors (TSCs) kitaev2001unpaired; kitaev2003fault; qi2011topological; leijnse2012introduction; beenakker2013search; aguado2017majorana. The MSs obey a non-Abelian exchange statistic, which makes them promising candidates for realizing robust qubits with potential applications in fault-tolerant quantum computing ivanov2001non; nayak2008non; alicea2011non; aasen2016milestones.
@@
Topological superconductivity (TS) can be engineered by using semiconductor nanowires with large spin-orbit coupling (SOC) and proximitized by s-wave superconductors PhysRevLett.105.177002; PhysRevB.82.214509; PhysRevLett.104.040502; PhysRevLett.105.077001; rokhinson2012fractional; PhysRevLett.109.227006; mourik2012signatures; das2012zero; deng2012anomalous; deng2016majorana; manna2020signature, proximitized systems exposed to magnetic textures PhysRevLett.109.236801; PhysRevB.85.020503; PhysRevLett.117.077002; matos2017tunable; PhysRevApplied.12.034048; PhysRevB.99.134505; PhysRevB.95.140504; desjardins2019synthetic; PhysRevB.104.174502; Marra2024:JAP; Kovalev2022:JAP, and magnetic chains on s-wave superconductors PhysRevB.84.195442; PhysRevB.85.144505; PhysRevB.88.155420; PhysRevB.88.020407; nadj2014observation; pawlak2016probing. Proximitized planar JJs have recently emerged as promising platforms for creating and manipulating MSs PhysRevLett.89.137007; Pientka2017:PRX; PhysRevLett.111.107007; PhysRevLett.118.107701; Setiawan2019:PRBb; Setiawan2019:PRBa; PhysRevB.89.195407; fornieri2019evidence; ren2019topological; PhysRevLett.126.036802; hart2014induced; PhysRevLett.125.086802; PhysRevB.103.L180505; lesser2021phase; zhou2022fusion; PhysRevB.99.035307; PhysRevB.96.205425; PhysRevB.99.214503; PhysRevLett.124.137001; PhysRevB.102.245403; PhysRevB.101.195435; PhysRevB.104.155428; salimian2021gate; PhysRevB.95.174515; PhysRevResearch.4.043087; PhysRevB.107.245304; Pekerten2024:PRB. In addition to the experimental advances in building such structures, proximitized planar JJs have also been shown to possess an enhanced parameter space supporting the topological superconducting state Pientka2017:PRX; Pekerten2024:APL.
@@
Magnetic and crystalline anisotropic effects have been predicted to appear in the Josephson junction with noncentrosymmetric materials CATSJJ; Pekerten2022:PRB; PhysRevB.56.892; PhysRevLett.105.097002. In particular, it has been shown that in the presence of SOC the Zeeman interaction yields a strong dependence of the system properties on the magnetic field direction. Furthermore, in systems with Rashba bychkov1984oscillatory and Dresselhaus dresselhaus1955spin SOCs the crystallographic orientation can affect the topological superconducting state, its robustness, and signatures Scharf2019:PRB; CATSJJ; Pekerten2022:PRB.
@@
Most previous investigations of TS in planar JJs have focused on end MSs, i.e., MSs that localize at the opposite ends of the junction with short localization lengths both along the junction ($\hat{y}$ direction, as shown in Fig. 1) and along the system edges perpendicular to the junction ($\hat{x}$ direction, as shown in Fig. 1) Pekerten2022:PRB; PhysRevLett.124.137001; CATSJJ; PhysRevB.107.035435; zhou2022fusion. However, theoretical evidence of the existence of edge-like MSs (i.e., MSs that are localized along the junction direction but extend along the edges perpendicular to the junction over distances significantly larger than the junction width) has been provided in previous works Sau2010:PRL; fornieri2019evidence; CATSJJ. The formation of MSs exhibiting anomalous multilocality in three-terminal Josephson junctions has also been proposed PhysRevB.110.L041110. Besides JJs, Majorana edge states can also emerge in hybrid superconductor/ferromagnet structures with helical magnetic textures wu2017:PRB. Moreover, two-dimensional structures, typically associated with quantum anomalous Hall systems coupled to superconductors, have been shown to support chiral Majorana edge states Lian2018:PNAS; menard2019:Nature; Beenakker2020:SP flowing around the sample edges in opposite directions and Majorana corner modes localized at vertices PhysRevB.97.205134; PhysRevB.109.115413; PhysRevB.109.205158.
@@ II Theoretical Model
Figure 1: (a) Schematic of a JJ consisting of a noncentrosymmetric semiconductor 2DEG (blue) in contact with two superconducting (S) leads (green). The $\hat{x}$ and $\hat{y}$ axes define the coordinate system in the junction’s reference frame. A top gate (not shown) over the normal (N) region can be used to modulate the Rashba SOC strength PhysRevLett.126.036802; mayer2020gate. (b) Relevant angles in the junction coordinate system: $\varphi_{B}$ defines the direction of the in-plane magnetic field ($\mathbf{B}$) with respect to the $\hat{x}$ axis, while $\theta_{c}$ determines the orientation of the junction reference frame with respect to the semiconductor’s [100] crystallographic axis.
@@ II.1 Topological charge
where $\rm Pf$$\{$...\} denotes the Pfaffian schnyder2008classification; ryu2010topological; ghosh2010non; Pfaffian. The topological charge determines whether a system belonging to the D class is in a trivial ($Q=1$) or topological (Q=1Q=-1) phase rieder2013reentrant; adagideli2014effects; pekerten2017disorder; stanescu2011majorana; stanescu2013majorana; pekerten2019fermion.
It is worth noting that under some conditions determined by the SOC field, the magnetic field direction, and the junction crystallographic orientation, symmetric junctions may effectively belong to the BDI class CATSJJ. In such cases, the topological phases are characterized by the $\mathbb{Z}$ topological invariant of the BDI class. Since the topological charge $Q$ is determined by the parity of the $\mathbb{Z}$ index, the topologically non-trivial regions of the BDI class consist of non-trivial D-class regions (composed of odd $\mathbb{Z}$ index subregions) enhanced by regions with even $\mathbb{Z}$ index Pientka2017:PRX; Setiawan2019:PRBb. We found that the topological gap in BDI class regions with multiple pairs of MSs is relatively small in the systems considered here. Therefore, our investigation focuses on regions that support only a single pair of MSs. The extent of these regions can be determined by examining how the topological charge $Q$ depends on the system parameters.
@@ II.2 Topological gap
Here $E_{0}$ and $E_{1}$ are the two lowest-energy states on the positive branch of the energy spectrum, respectively. Due to finite-size effects, the MSs localized at opposite ends (or edges) may overlap, so their energy ($\pm E_{0}$) may slightly deviate from zero. Note that Eq. (8) can only be interpreted as the topological gap when the system is in the TS state.
@@
The magnitude of $\Delta_{top}$ has been shown to strongly depend on the junction’s crystallographic orientation $(\theta_{c})$, the spin-orbit angle $(\theta_{so})$ and the in-plane magnetic field orientation $(\varphi_{B})$, being optimal when the following relation is fulfilled CATSJJ; Pekerten2022:PRB,
@@
As briefly discussed in the Introduction, end-like MSs are localized at the opposite ends of the junction channel, with short localization lengths both along the junction and along the system’s edges perpendicular to it. In contrast, edge-like MSs are strongly localized along the junction direction but extend over distances significantly larger than the junction width along the system’s edges perpendicular to the junction. We want to emphasize that the edge-like MSs examined here differ from both chiral Majorana states commonly associated with quantum anomalous Hall systems coupled to superconductors Lian2018:PNAS; menard2019:Nature; Beenakker2020:SP and unidirectional Majorana edge states in noncentrosymmetric superconductors Daido2017:PRB.
To understand the origin of the extended nature of edge-like MSs, consider a simplified model of a junction sufficiently long such that the overlap between Majorana states at opposite edges is negligible. In this scenario, we can focus solely on one Majorana state, for instance, the one localized at the bottom edge ($y=0$). Assuming the S regions are infinitely wide ($W_{S}\to\infty$), the scattering states in the left S region can be expressed as CATSJJ; Scharf2019:PRB,
@@
where $\kappa$ describes the localization of the Majorana state along the junction $y$-direction, the subindexes $e$ and $h$ refer to electron-like and hole-like states, respectively, $s=\pm$ characterizes the spin, and $q_{e,s}$ ($q_{h,s}$) is the wave vector of the electron-like (hole-like) state with spinor $\chi_{e,s}$ ($\chi_{h,s}$) and spin $s$. In Eq. (10), we have omitted the oscillations of the MSs along the junction direction, as they become negligibly small for sufficiently long junctions.
@@
The overall localization length of the Majorana state along the edge is given by the largest of the $l_{j,s}$, which may become significantly larger than the junction width for wave vectors with a sufficiently small imaginary part, leading to the formation of edge-like MSs. At $\phi=0$, if at least one wave vector is purely real, the corresponding scattering mode represents a propagating wave with infinite localization length. In this particular case, the edge-like MS propagates without decay over the entire edge. Conversely, end-like MSs appear when all wave vectors are complex.
@@
are satisfied, all coefficients of the quartic equation become real. This allows for purely real wave vectors, enabling the emergence of propagating edge-like MSs if the system parameters are appropriately tuned. Note that Eq. (14) is a necessary, though not sufficient, condition for the emergence of propagating edge-like MSs in junctions with infinitely wide S regions.
@@
Propagating edge-like MSs with infinite localization length cannot be practically realized, as their extension is ultimately limited by the finite size of the S region in realistic JJs. Therefore, we focus our numerical investigation on edge-like MSs with negligibly small decay over the length of the S regions, which, in the numerical simulations, are set to be much larger than the junction width. To ensure that the edge-like MSs extend across the entire S regions of the finite-size junctions considered, we impose the necessary constraints derived from the asymptotic analysis in the limit $W_{S}\rightarrow\infty$. Consequently, all configurations in the numerical analysis presented in Sec. III are chosen such that Eq. (14) and the condition for optimal topological gap [Eq. (9) are simultaneously satisfied.
@@
$\tilde{\Delta}=\dfrac{(1-Q)\Delta_{top}}{2\Delta_{0}}\,{\rm sgn}(\eta)$ (15)
where $\eta$ is a parameter that characterizes the nature of the MSs, taking positive values for edge-like MSs and negative values for end-like MSs.
@@
[…]
@@
For the junctions studied here, the probability density of edge-like MSs displays an oscillatory behavior along the edge, with multiple local maxima of comparable amplitudes. In contrast, the probability density oscillations of end-like MSs decay rapidly, featuring a large maximum at the middle of the junction edge, followed by a few significantly smaller maxima. Therefore, the localization nature of the MSs can be captured by the parametrization $\eta=n-n_{0}-1/2$, where $n$ is the number of probability density local maxima with comparable amplitudes. Since the typical number of local maxima for end-like MSs is smaller than 3, we set $n_{0}=3$ in the numerical simulations of Eq.(16) discussed below. For higher chemical potentials and/or larger $W_{N}$ and $W_{S}$ one may need to increase $n_{0}$ accordingly. Note that other equivalent definitions of $\eta$ are also possible. For instance, one could define $\eta$ to be positive when the MS localization length is much larger than the edge length $W$, and negative otherwise. However, the two approaches are equivalent, since local maxima of the probability density extend along the entire edge with comparable amplitudes only when the localization length significantly exceeds $W$.
@@ III Results
We consider two types of junctions: (i) Al/HgTe Josephson junctions (JJs), where Rashba spin-orbit coupling (SOC) is the dominant effect, and (ii) Al/InSb JJs, where both Rashba and Dresselhaus SOC may be significant. The system parameters used in our calculations are provided in Appendix A. The numerical simulations were carried out by discretizing Eq. (1) on a mesh with a lattice constant of $a=10$ nm. Using the finite-difference approximation, we constructed the tight-binding (TB) form of the BdG Hamiltonian with the Kwant package groth2014kwant. The energy spectrum and wave functions were obtained by numerically diagonalizing the TB BdG Hamiltonian on a finite lattice. The energy spectrum was then used to compute the topological gap [Eq. (8)], the phase-dependent part of the free energy [Eq. (16)], and the ground-state phase [Eq. (16)]. The topological charge [see Eq. (7)] was computed by using the system TB BdG Hamiltonian with imposed translational invariance along the junction direction.